# Kernel Banzhaf: A Fast and Robust Estimator for Banzhaf Values

### v1

@@ Appendix A Theoretical Guarantees
$\|\mathbf{A\tilde{\bm{\phi}}-A\bm{\phi}}\|_{2}^{2}=\|\tilde{\mathbf{y}}-\mathbf{y}\|_{2}^{2}\leq 2\|\mathbf{U^{\top}S^{\top}SU}(\tilde{\mathbf{y}}-\mathbf{y})\|_{2}\leq\epsilon\|\mathbf{Uy-b}\|_{2}=\epsilon\|\mathbf{A}\bm{\phi}-\mathbf{b}\|_{2}^{2}$ (39)
@@
Appendix B Implementation Details
Feature perturbation and exact Banzhaf value calculation
There are generally two approaches to handling removed features in feature perturbation for general set functions, as discussed in Chen et al. (2020) and Kumar et al. (2020). Given an explicand $x$ and a subset of features $S$, define $\mathbf{x}^{S}$ as the observation where $\mathbf{x}^{S}_{i}=\mathbf{x}^{e}_{i}$ if feature $i\in S$ and, otherwise, $\mathbf{x}^{\bar{S}}$ is sampled from a distribution. The first method involves sampling from the conditional distribution of the removed features, where replacement values for the absent features are sampled according to $\mathbf{x}^{\bar{S}}\sim p(\mathbf{x}^{\bar{S}}|\mathbf{x}^{S})$. This approach, while precise, is computationally expensive. Alternatively, the marginal distribution can be used where the observed features $\mathbf{x}^{S}$ are ignored, and replacement values are sampled according to $\mathbf{x}^{\bar{S}}\sim p(\mathbf{x}^{\bar{S}})$. Due to its lower computational complexity, we adopt the latter approach.
For implementation, distinct feature perturbation methods are applied depending on the set function, i.e. model type. For tree-based models, as we need to use the tree-based algorithm for calculating the ground truth Banzhaf values to measure errors, we utilize a method aligned with Algorithm 1 from Karczmarz et al. (2022), which computes predictions using partial features. Specifically, during tree traversal, if feature $i\in S$, we proceed according to the threshold to select the child node; if $i\notin S$, we traverse both children and compute a weighted average of the predictions, effectively nullifying the influence of features not in $S$ without any feature value replacement.
For neural network models, instead of using fixed baseline values for the removed features, we compute the average of the model’s predictions using replacement values randomly sampled from 50 baseline points, different from the explicand. The explicand $x^{e}$ is repeated 50 times, with the non-selected features replaced by values from these baseline points, and the average of $M(x^{S})$ is taken to estimate the impact of marginalizing out the non-selected features. To calculate ground truth Banzhaf values, we evaluate all $2^{n}$ subsets of features in this way.
Sample size for Monte Carlo
Given a sample size of $n$, the MSR and Kernel Banzhaf estimators effectively use $n$ samples to estimate Banzhaf values for all features concurrently. In contrast, the Monte Carlo method estimates each feature independently and requires a fair allocation of the total $m$ samples. To achieve this, we implement a loop where for each iteration $i=1,\ldots,m$, we use $i\mod n+1$ samples when $i$ is divisible by $n$, and $i\mod n$ samples otherwise. This approach ensures that the Monte Carlo method also utilizes exactly $m$ samples, maintaining consistency in sample usage across different estimators.
@@ Appendix C Datasets and Models
Appendix D Neural Network-Based Model Explanations
Figure 6: $\ell_{2}$-norm error by sample size for Banzhaf estimators in explaining neural network models. This figure compares the $\ell_{2}$-norm errors of Kernel Banzhaf (including an ablated version without paired sampling), MC, and MSR across increasing sample sizes on four small datasets. The results highlight the robust performance and generalizability of Kernel Banzhaf across various model types.
@@ Appendix D Neural Network-Based Model Explanations
Appendix E Complementary Banzhaf Estimator Experiments
Figure 7: Relative objective error based on $\|\mathbf{Ax-b}\|_{2}$ plotted against sample size (top row) and noise level (bottom row) for comparing Banzhaf estimators across four small datasets. Kernel Banzhaf consistently outperforms other estimators, demonstrating enhanced accuracy and robustness as sample sizes and noise levels vary. This performance highlights its efficacy in solving the linear regression problem where the solutions represent Banzhaf values.
Figure 8: Computational time required for Banzhaf value estimation across varying sample sizes for eight datasets. Notably, Kernel Banzhaf (both with and without paired sampling) and MSR demonstrate comparable computational efficiencies, while MC consistently requires approximately twice the computation time, particularly evident when sample size is small.
Relative Objective Error
@@ Appendix E Complementary Banzhaf Estimator Experiments
To further demonstrate the superior performance of Kernel Banzhaf, we evaluated the Banzhaf estimators against exact Banzhaf values using relative objective error. Building on the framework where Banzhaf values solve a linear regression problem (as established in Theorem 3.2), we aim to minimize the objective function $\|\mathbf{Ax-b}\|_{2}$. Here, $\bm{\phi}$ represents the vector of exact Banzhaf values for each feature, and $\hat{\bm{\phi}}$ represents the estimations from different estimators. We evaluate the differences between $\|\mathbf{A\bm{\phi}-b}\|_{2}$ and $\|\mathbf{A\hat{\bm{\phi}}-b}\|_{2}$, because although the optimal linear regression solution corresponds to the Banzhaf values, the optimal objective error $\|\mathbf{A\bm{\phi}-b}\|_{2}$ is not necessarily zero. Thus, we assess the relative objective error. The comparative analysis, shown in Figure 7, reveals that the plots of this error with increasing sample size and noise level are analogous to those using $\ell_{2}$-norm error, where Kernel Banzhaf consistently surpasses all other estimators in both efficacy and robustness. These experiments underscore the effectiveness of our Kernel Banzhaf algorithm in accurately estimating Banzhaf values and solving the corresponding linear regression challenge.
Time Complexity
@@
We further evaluated the computational efficiency of the MC, MSR, and Kernel Banzhaf estimators by measuring the exact time required to estimate Banzhaf values, as depicted in Figure 8. All experiments were conducted on a Lenovo SD650 with 128 GB of RAM, using only one thread for computation. Our analysis reveals that both MSR and Kernel Banzhaf (with and without paired sampling) exhibit comparable computational times across all datasets and sample sizes. In contrast, the MC method consistently requires approximately twice the time of the other estimators. This discrepancy arises primarily because the most time-consuming operation involves computing the set function output $v(S)$ for a given sample $S$ of players. Kernel Banzhaf and MSR efficiently leverage a single calculation of $v(S)$ per sample for all players. In contrast, MC needs to compute both $v(S\cup i)$ and $v(S)$ to determine the marginal contribution for each feature $i$ individually, thus doubling the computation time compared to the other estimators.
Appendix F Complementary Kernel Banzhaf vs. KernelSHAP Experiments
Figure 9: Normalized $\ell_{2}$-norm error (i.e. $\|\hat{\bm{\phi}}-\bm{\phi}\|_{2}^{2}/\|\bm{\phi}\|_{2}^{2}$) by sample size for Kernel Banzhaf and KernelSHAP estimators. This set of plots compares the $\ell_{2}$-norm error across increasing sample sizes for Kernel Banzhaf, Leverage SHAP, and Optimized KernelSHAP across eight datasets. The plots aim to evaluate the efficiency and accuracy of Kernel Banzhaf (the best Banzhaf estimator) against advanced KernelSHAP methods under varying sample sizes. It demonstrates that Kernel Banzhaf outperforms KernelSHAP algorithms for larger feature sets.
In this section, we assess the performance of Kernel Banzhaf against state-of-the-art KernelSHAP algorithms by varying sample sizes. Kernel Banzhaf, as shown in Figure 9, demonstrates superior performance on larger datasets (bottom row). For smaller datasets, the three algorithms appear comparable at first glance, since KernelSHAP are using sampling without replacement to increase the diversity of the samples; however, the smaller shaded areas for Kernel Banzhaf, representing the 25% and 75% percentiles, indicate significantly greater stability compared to KernelSHAP algorithms, which highlight Kernel Banzhaf’s enhanced stability.
Appendix G Kernel Banzhaf with Sampling without Replacement

### v2

@@ Appendix A Kernel Banzhaf Approximation Guarantees
$\|\mathbf{A\tilde{\boldsymbol{\phi}}-A\boldsymbol{\phi}}\|_{2}^{2}=\|\tilde{\mathbf{y}}-\mathbf{y}\|_{2}^{2}\leq 2\|\mathbf{U^{\top}S^{\top}SU}(\tilde{\mathbf{y}}-\mathbf{y})\|_{2}\leq\epsilon\|\mathbf{Uy-b}\|_{2}=\epsilon\|\mathbf{A}\boldsymbol{\phi}-\mathbf{b}\|_{2}^{2}$ (37)
@@
Appendix B Extension to Probabilistic Values
In this section, we consider extending Kernel Banzhaf to probabilistic values. We will first establish how the optimal solution connects to the probabilistic values. Next, we will prove guarantees on the approximate solution selected via leverage score sampling.
In the remainder of this section, let $\mathbf{A}\in\mathbb{R}^{2^{n}\times n}$ such that
$[\mathbf{A}]_{S,i}=\begin{cases}\frac{p_{|S|-1}}{a_{n}}&\textrm{if }i\in S\\ -\frac{p_{|S|}}{a_{n}}&\textrm{if }i\notin S\\ \end{cases}$ (38)
where $i\in[n]$ and we use $S\subseteq[n]$ as an index for ease of notation. Let $\mathbf{b}\in\mathbb{R}^{2^{n}}$ be the vector representation of the set function $v$ i.e., $b_{S}=v(S)$.
B.1 Equivalence to Probabilistic Values
See 3.4
Proof of Lemma 3.4.
Our proof strategy is to explicitly compute the optimal solution
$\mathbf{x}^{*}=(\mathbf{A}^{\top}\mathbf{A})^{-1}\mathbf{A}^{\top}\mathbf{b}.$ (39)
We begin with $\mathbf{A}^{\top}\mathbf{A}$. In the computation, we will repeatedly use the symmetry of $\mathbf{A}^{\top}\mathbf{A}$; namely, all the diagonal entries are the same and all the non-diagonal entries are the same. The diagonal entries for $i\in[n]$ are given by
$a_{n}^{2}[\mathbf{A}^{\top}\mathbf{A}]_{i,i}$ $=\sum_{S\subseteq[n]:i\in S}p_{|S|-1}^{2}+\sum_{S\subseteq[n]:i\notin S}p_{|S|}^{2}=\sum_{\ell=1}^{n}\binom{n-1}{\ell-1}p_{\ell-1}^{2}+\sum_{\ell=0}^{n-1}\binom{n-1}{\ell}p_{\ell}^{2}$ (40)
$=2\sum_{\ell=0}^{n-1}\binom{n-1}{\ell}p_{\ell}^{2}$ (41)
where the last equality follows by a change of variable. Similarly, the off-diagonal entries for $i\neq j$ are given by
$a_{n}^{2}[\mathbf{A}^{\top}\mathbf{A}]_{i,j}$ $=\sum_{S\subseteq[n]:i,j\in S}p_{|S|-1}^{2}-2\sum_{S\subseteq[n]:i\in S,j\notin S}p_{|S|-1}p_{|S|}+\sum_{S\subseteq[n]:i,j\notin S}p_{|S|}^{2}$ (42)
$=\sum_{\ell=2}^{n}\binom{n-2}{\ell-2}p_{\ell-1}^{2}-2\sum_{\ell=1}^{n-1}\binom{n-2}{\ell-1}p_{\ell-1}p_{\ell}+\sum_{\ell=0}^{n-2}\binom{n-2}{\ell}p_{\ell}^{2}$ (43)
$=\sum_{\ell=1}^{n-1}\binom{n-2}{\ell-1}\left(p_{\ell}-p_{\ell-1}\right)^{2}=a_{n}b_{n}.$ (44)
By the definition of $a_{n}$, we have
$a_{n}$ $=2\sum_{\ell=0}^{n-1}\binom{n-1}{\ell}p_{\ell}^{2}-\sum_{\ell=1}^{n-1}\binom{n-2}{\ell-1}\left(p_{\ell}-p_{\ell-1}\right)^{2}=a_{n}^{2}[\mathbf{A}^{\top}\mathbf{A}]_{i,i}-a_{n}^{2}[\mathbf{A}^{\top}\mathbf{A}]_{i,j}.$ (45)
It follows that $[\mathbf{A}^{\top}\mathbf{A}]_{i,i}-[\mathbf{A}^{\top}\mathbf{A}]_{i,j}=\frac{1}{a_{n}}.$ Then
$\mathbf{A}^{\top}\mathbf{A}$ $=([\mathbf{A}^{\top}\mathbf{A}]_{i,i}-[\mathbf{A}^{\top}\mathbf{A}]_{i,j})\mathbf{I}+[\mathbf{A}^{\top}\mathbf{A}]_{i,j}\mathbf{1}\mathbf{1}^{\top}=\frac{1}{a_{n}}(\mathbf{I}+b_{n}\mathbf{1}\mathbf{1}^{\top}).$ (46)
By the Sherman-Morrison formula, $(\mathbf{I}+b_{n}\mathbf{1}\mathbf{1}^{\top})^{-1}=\mathbf{I}-\frac{b_{n}}{1+nb_{n}}\mathbf{1}\mathbf{1}^{\top}$. Then
$(\mathbf{A}^{\top}\mathbf{A})^{-1}=a_{n}\left(\mathbf{I}-\frac{b_{n}}{1+nb_{n}}\mathbf{1}\mathbf{1}^{\top}\right).$ (47)
Next, we compute $\mathbf{A}^{\top}\mathbf{v}$. The $i$th entry of this $n\times 1$ vector is given by
$a_{n}[\mathbf{A}^{\top}\mathbf{v}]_{i}$ $=\sum_{S\subseteq[n]:i\in S}p_{|S|-1}v(S)-\sum_{S\subseteq[n]:i\notin S}p_{|S|}v(S)=\sum_{S\subseteq[n]\setminus\{i\}}p_{|S|}(v(S\cup\{i\})-v(S))=\phi_{i}.$ (48)
Therefore $\mathbf{A}^{\top}\mathbf{v}=\frac{1}{a_{n}}\boldsymbol{\phi}$. Finally,
$\mathbf{x}^{*}$ $=(\mathbf{A}^{\top}\mathbf{A})^{-1}\mathbf{A}^{\top}\mathbf{b}=a_{n}\left(\mathbf{I}-\frac{b_{n}}{1+nb_{n}}\mathbf{1}\mathbf{1}^{\top}\right)\frac{1}{a_{n}}\boldsymbol{\phi}$
and, taking the inverse,
$\phi$ $=\left(\mathbf{I}-\frac{b_{n}}{1+nb_{n}}\mathbf{11^{\top}}\right)^{-1}\mathbf{x}^{*}=(\mathbf{I}+b_{n}\mathbf{11^{\top}})\mathbf{x}^{*}.$ (49)
B.2 Approximation Guarantees
See 3.5
We prove Theorem 3.5 by applying Theorem A.1 to the generalized regression problem. First, we compute the leverage scores of the generalized regression problem.
Lemma B.1.
Let $S\subseteq[n]$ be an index. The corresponding leverage score of the generalized regression problem is
$\ell_{S}$ $=\frac{1}{a_{n}}\biggl{[}|S|p_{|S|-1}^{2}+(n-|S|)p_{|S|}^{2}-\frac{b_{n}}{1+nb_{n}}\left(|S|p_{|S|-1}-(n-|S|)p_{|S|})^{2}\right)\biggr{]}.$ (50)
Proof of Lemma B.1.
Using the definition of leverage scores and Equation 47, we have
$\ell_{S}$ $=[\mathbf{A}]_{S}^{\top}(\mathbf{A}^{\top}\mathbf{A})^{-1}[\mathbf{A}]_{S}=a_{n}[\mathbf{A}]_{S}^{\top}\left(\mathbf{I}-\frac{b_{n}}{1+nb_{n}}\mathbf{1}\mathbf{1}^{\top}\right)[\mathbf{A}]_{S}=a_{n}\left(\|[\mathbf{A}]_{S}\|_{2}^{2}-\frac{b_{n}}{1+nb_{n}}\|[\mathbf{A}]_{S}\|_{1}^{2}\right).$ (51)
By the definition of $\mathbf{A}$, the $\ell_{2}$-norm squared is given by
$\|[\mathbf{A}]_{S}\|_{2}^{2}$ $=\frac{1}{a_{n}^{2}}\left(\sum_{i\in S}p_{|S|-1}^{2}+\sum_{i\notin S}p_{|S|}^{2}\right)=\frac{1}{a_{n}^{2}}\left(|S|p_{|S|-1}^{2}+(n-|S|)p_{|S|}^{2}\right)$ (52)
while the $\ell_{1}$-norm squared is given by
$\|[\mathbf{A}]_{S}\|_{1}^{2}$ $=\frac{1}{a_{n}^{2}}\left(\sum_{i\in S}p_{|S|-1}-\sum_{i\notin S}p_{|S|}\right)^{2}=\frac{1}{a_{n}^{2}}\left(|S|p_{|S|-1}-(n-|S|)p_{|S|}\right)^{2}.$ (53)
The lemma statement follows. ∎
Next, we prove Theorem 3.5.
Proof of Theorem 3.5.
Let $\gamma=\|\mathbf{Ax}^{*}-\mathbf{b}\|_{2}^{2}/\|\mathbf{Ax}^{*}\|_{2}^{2}$ and suppose there are $m=O(n\log(n/\delta)+n\log(1/\delta)/\epsilon)$ samples. By the standard leverage score sampling guarantee in Theorem A.1 (see e.g., the first few lines of the proof of Corollary 3.3 for the step-by-step process), we have
$\|\mathbf{A}\tilde{\mathbf{x}}-\mathbf{A}\mathbf{x}^{*}\|_{2}^{2}\leq\epsilon\|\mathbf{A}\mathbf{x}^{*}-\mathbf{b}\|_{2}^{2}=\epsilon\gamma\|\mathbf{A}\mathbf{x}^{*}\|_{2}^{2}$ (54)
with probability $1-\delta$. For the data matrix $\mathbf{A}$ in the unconstrained regression problem, we have
$\|\mathbf{Ax}\|_{2}^{2}=\mathbf{x^{\top}A^{\top}Ax}=\frac{1}{a_{n}}\mathbf{x}^{\top}(\mathbf{I}+b_{n}\mathbf{11}^{\top})\mathbf{x}=\frac{1}{a_{n}}\left[\|\mathbf{x}\|_{2}^{2}+b_{n}(\mathbf{1^{\top}x})^{2}\right]$ (55)
where the second equality follows by Equation 46.
Then the leverage score sampling guarantee implies
$\frac{1}{a_{n}}\left[\|\mathbf{x}^{*}-\tilde{\mathbf{x}}\|_{2}^{2}+b_{n}(\mathbf{1}^{\top}(\mathbf{x}^{*}-\tilde{\mathbf{x}}))^{2}\right]\leq\epsilon\gamma\frac{1}{a_{n}}\left[\|\mathbf{x}^{*}\|_{2}^{2}+b_{n}(\mathbf{1}^{\top}\mathbf{x}^{*})^{2}\right]$ (56)
so
$\|\mathbf{x}^{*}-\tilde{\mathbf{x}}\|_{2}^{2}\leq\epsilon\gamma\left(\|\mathbf{x}^{*}\|_{2}^{2}+b_{n}(\mathbf{1}^{\top}\mathbf{x}^{*})^{2}\right)\leq\epsilon\gamma(1+nb_{n})\|\mathbf{x}^{*}\|_{2}^{2}$ (57)
where the last inequality follows because $(\mathbf{1}^{\top}\mathbf{x}^{*})^{2}\leq\|\mathbf{1}\|_{2}^{2}\|\mathbf{x}^{*}\|_{2}^{2}=n\|\mathbf{x}^{*}\|_{2}^{2}$ by Cauchy-Schwarz.
Since $b_{n}$ is non-negative, observe that $\|\mathbf{I}+b_{n}\mathbf{11}^{\top}\|_{2}=1+nb_{n}$ and $\|(\mathbf{I}+b_{n}\mathbf{11}^{\top})^{-1}\|_{2}=1$. It follows that
$\|\mathbf{x}^{*}\|_{2}^{2}=\|(\mathbf{I}+b_{n}\mathbf{11}^{\top})^{-1}(\mathbf{I}+b_{n}\mathbf{11}^{\top})\mathbf{x}^{*}\|_{2}^{2}=\|(\mathbf{I}+b_{n}\mathbf{11}^{\top})^{-1}\boldsymbol{\phi}\|_{2}^{2}\leq\|(\mathbf{I}+b_{n}\mathbf{11}^{\top})^{-1}\|_{2}^{2}\cdot\|\boldsymbol{\phi}\|_{2}^{2}=\|\boldsymbol{\phi}\|_{2}^{2}.$ (58)
Then
$\|\boldsymbol{\phi}-\tilde{\boldsymbol{\phi}}\|_{2}^{2}$ $=\|(\mathbf{I}+b_{n}\mathbf{11}^{\top})(\mathbf{x}^{*}-\tilde{\mathbf{x}})\|_{2}^{2}$ (59)
$\leq\|\mathbf{I}+b_{n}\mathbf{11}^{\top}\|_{2}^{2}\cdot\|\mathbf{x}^{*}-\tilde{\mathbf{x}}\|_{2}^{2}$ (60)
$\leq\epsilon\gamma(1+nb_{n})^{3}\|\mathbf{x}^{*}\|_{2}^{2}$ (61)
$\leq\epsilon\gamma(1+nb_{n})^{3}\|\boldsymbol{\phi}\|_{2}^{2}$ (62)
@@ Appendix C Datasets and Models
There are generally two approaches to handling removed features in feature perturbation for general set functions, as discussed in Chen et al. (2020) and Kumar et al. (2020). Given an explicand $x$ and a subset of features $S$, define $\mathbf{x}^{S}$ as the observation where $\mathbf{x}^{S}_{i}=\mathbf{x}_{i}$ if feature $i\in S$ and, otherwise, $\mathbf{x}^{S}_{i}$ is sampled from one of two distributions. The first method involves sampling from the conditional distribution of the removed features. This approach, while precise, is computationally expensive. Alternatively, the marginal distribution can be used where the observed features $\mathbf{x}^{S}_{i}$ for $i\in S$ are ignored. Due to its lower computational complexity, we adopt the latter approach.
For tree-based models, as we need to use the tree-based algorithm for calculating the ground truth Banzhaf values to measure errors, we utilize a method aligned with Algorithm 1 from Karczmarz et al. (2022), which computes predictions using partial features. Specifically, during tree traversal, if feature $i\in S$, we proceed according to the threshold to select the child node; if $i\notin S$, we traverse both children and compute a weighted average of the predictions, effectively nullifying the influence of features not in $S$ without any feature value replacement. We use XGBoost (Chen & Guestrin, 2016) for our tree-based models, which makes it possible to apply the tree-based algorithm of (Karczmarz et al., 2022) to compute exact Banzhaf values for $n\geq 30$.
For neural network models, instead of using fixed baseline values for the removed features, we compute the average of the model’s predictions using replacement values randomly sampled from 50 baseline points, different from the explicand. For each explicand $\mathbf{x}$, the non-selected features in $\mathbf{x}^{S}$ are replaced by values from baseline points, and the average of $M(\mathbf{x}^{S})$ is taken to estimate the impact of marginalizing out the non-selected features. To calculate ground truth Banzhaf values, we evaluate all $2^{n}$ subsets of features in this way.
Appendix D Neural Network Models
@@
Figure 5: This figure compares the relative squared $\ell_{2}$-norm errors of Kernel Banzhaf (including an ablated version without paired sampling), MC, and MSR across increasing sample sizes on four small datasets in explaining neural network models. The results highlight the robust performance and generalizability of Kernel Banzhaf across various model types.
Appendix E Alternative Error Metric: Regression Objective Error
@@
To further demonstrate the superior performance of Kernel Banzhaf, we evaluated the Banzhaf estimators against exact Banzhaf values using relative objective error in Figure 6. Building on the framework where Banzhaf values solve a linear regression problem (as established in Theorem 3.1), we aim to minimize the objective function $\|\mathbf{Ax-b}\|_{2}$. Here, $\boldsymbol{\phi}$ represents the vector of exact Banzhaf values for each feature, and $\tilde{\boldsymbol{\phi}}$ represents the estimations from different estimators. We evaluate the differences between $\|\mathbf{A\boldsymbol{\phi}-b}\|_{2}$ and $\|\mathbf{A\tilde{\boldsymbol{\phi}}-b}\|_{2}$, because although the optimal linear regression solution corresponds to the Banzhaf values, the optimal objective error $\|\mathbf{A\boldsymbol{\phi}-b}\|_{2}$ is not necessarily zero. Thus, we assess the relative objective error. The comparative analysis, shown in Figure 6, reveals that the plots of this error with increasing sample size and noise level are analogous to those using relative squared $\ell_{2}$-norm error, where Kernel Banzhaf consistently surpasses all other estimators in both efficacy and robustness. These experiments underscore the effectiveness of our Kernel Banzhaf algorithm in accurately estimating Banzhaf values and solving the corresponding linear regression challenge.
Figure 6: Relative objective error $\frac{\|\mathbf{A\tilde{\boldsymbol{\phi}}-b}\|_{2}-\|\mathbf{A\boldsymbol{\phi}-b}\|_{2}}{\|\mathbf{A\boldsymbol{\phi}-b}\|_{2}}$ plotted against sample size (top row) and noise level (bottom row) for comparing Banzhaf estimators across four small datasets. Kernel Banzhaf consistently outperforms other estimators, demonstrating enhanced accuracy and robustness as sample sizes and noise levels vary. This performance highlights its efficacy in solving the linear regression problem where the solutions represent Banzhaf values.
Appendix F Time Complexity
@@
We further evaluated the computational efficiency of the MC, MSR, and Kernel Banzhaf estimators by measuring the exact time required to estimate Banzhaf values, as depicted in Figure 7. All experiments were conducted on a Lenovo SD650 with 128 GB of RAM, using only one thread for computation. Figure 7 reveals that all four estimators exhibit comparable computational times across all datasets and sample sizes, since evaluating the set function $v$ dominates the estimator’s time complexity.
Figure 7: Computational time required for Banzhaf value estimation across varying sample sizes for eight datasets. All estimators demonstrate comparable computational efficiencies primarily driven by the time required for model prediction.
Appendix G Ablation: Kernel Banzhaf With and Without Replacement
In order to ablate its sampling technique, we compare Kernel Banzhaf with and without replacement. Our empirical results, as illustrated in Figure 8, indicate that sampling without replacement does not yield improvements unless $m\geq 2^{n}$, in which case we can read the whole set function $v$ and exactly compute the Banzhaf values. We suspect the reason that Kernel Banzhaf does not benefit from sampling without replacement is because it uniformly samples subsets to solve the regression problem.